# post-last exit time process and its application to Loss-Given-Default Distribution

### v3

@@ 3.4. Model Check
Figure 4. Probability density function of (a) $K^{B}(\xi)$ in (3.7) and (b) total debt LGD $K^{D}(\xi)$ in (3.18) based on the estimated parameters
Table 4. Model-implied 5-year CDS spreads based on the distribution of $K^{D}(\xi)$, average LGD ($K^{D}(\xi)$), and their ratios (spread per $1\%$ LGD, denoted by $\hat{\rho}$) for each $w$ (up to 4 decimal points). Note that $\widehat{w}=70.1037\%$ and the quoted spread per $1\%$ LGD is $\rho_{\text{quote}}:=69.34/60=1.1557$bps.
w (%) 68 69 70.1037 71 72
CDS Spread (bps) 57.0172 57.4357 57.8976 58.2727 58.6912
Average LGD (%) 50.8359 51.2084 51.6195 51.9533 52.3258
$\widehat{\rho}$ 1.1216 1.1216 1.1216 1.1216 1.1217
@@
A.1. Proof of Lemma 2.1
@@
A.2. Detailed comments on Step 1 in Section 3.3
@@
We estimate the parameters of the asset process in (3.2) using the data of the firm’s equity $E$ and debt $B$ available up to the current time. The debt process $B$ represents a certain amount of debt to be repaid and we let it be the sum of short-term debt and a half of long-term debt (as in Moody’s KMV approach) since empirical evidence has demonstrated that such level is an appropriate default threshold for the firm’s assets. We use the option-theoretic approach of Merton (1974) where the value of equity represents the value of the European call option written on the firm’s assets with the strike price equal to the future value of debt. We estimate the asset parameters using the maximum likelihood method of Duan (1994), Duan (2000), and Lehar (2005). We assume that debt $B$ grows at the risk-free rate $r$. Assuming that debt $B$ is exogenous, the log-likelihood function to be maximized is given by
@@
A.3. Literature review of the study of loss-given default (LGD)
There exists a large body of literature regarding the loss-given-default (hereafter LGD) distribution because the information about this distribution is important for credit risk management. One example of its use is the estimation of loan-loss reserves in the banking industry. The LGD distribution should reflect credit market conditions as well as company-specific financial conditions. However, as Doshi et al. (2018) and many other studies point out, it is a standard practice to assume, irrespective of company names, a constant recovery rate (approximately 40%) of the debt upon default. This $60\%$ loss rate is arbitrarily determined, possibly from the empirical distribution. Since the market data is usually quoted based on the predetermined constant loss rate (60%) for all firms, it is hard to estimate the LGD distribution by just observing the market information such as CDS spreads, default time distribution, and default probability. In contrast to the standard assumption of the constant loss rate, the empirical literature has documented the evidence of time-varying realized loss rates. Hence we cannot overemphasize the effectiveness of a model that provides the LGD distribution implied in the credit market and company-specific financial conditions.
As the empirical literature has documented the evidence of time-varying realized loss rates, it is important to build a stochastic model for LGD. Gambetti et al. (2019) analyzes determinants of recovery rate distributions and finds economic uncertainty to be the most important systematic determinant of the mean and dispersion of the recovery rate distribution. For their analysis, the authors use post-default bond prices of 1831 American corporate defaults during 1990-2013. Even though recovery rates vary with a firm’s idiosyncratic factors, Gambetti et al. (2019) states that the impact of systematic factors related to economic cycles should not be underestimated. For the summary of studies related to cross-sectional and time variation of recovery rates, we refer the reader to Gambetti et al. (2019).
Finally, Cohen and Costanzino (2017) is an example of a study that extends a structural credit risk model and incorporates stochastic dynamics of the recovery rate. In contrast to our approach which does not introduce an additional recovery risk factor and is based solely on the dynamics of the leverage process, Cohen and Costanzino (2017) models asset and recovery processes separately as correlated geometric Brownian motions. In their model, the asset risk driver serves as a default trigger and the recovery risk driver determines the amount recovered upon default. The authors explicitly compute the prices of bonds and CDS under this framework. See also Kijima et al. (2009).

### v4

@@ 5.3. Analysis
Figure 1. Probability density function of total debt LGD $K^{D}(\xi)$ in (4.13) based on estimated parameters
Figure 2. Probability density function ${\rm d}\mathbb{P}_{y^{*}}(\lambda_{\alpha^{*}}\leq t)/{\rm d}t$, with $y^{*}=\frac{1}{\sigma}\ln(Y_{0})$, of the last passage time $\lambda_{\alpha^{*}}$ based on estimated parameters (including $\alpha$ and the starting point $Y_{0}$)
Looking at Figure 3 again, we see that the shape of the density for ‘LEN’ and ‘KSS’ is different from that of the remaining companies (all of which have a similar shape). The same pattern in observed in Figure 2. For these two companies, the densities for both $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c}$ are concentrated either within much narrower range (‘LEN’) or at much lower level (‘KSS’), compared to other companies, indicating high probability of the asset-to-debt ratio process not remaining above $\alpha^{*}$ for much longer. Based on Table 2, ‘LEN’ has the lowest drift $M$ and the highest expected value of $K^{D}(\xi)$ while ‘KSS’ has the highest 5Y PD. Thus, Figures 2 and 3 accurately capture high credit risk. Let us compare Figures 2 and 3 for all the companies; more specifically, Figure 2 and the density with $c=0$ in Figure 3. We see that densities of $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c}$ with $c=0$ put relatively similar masses in the same regions, which means that the process $\frac{1}{\sigma}\ln(Y)$ spends most of the time above level $\alpha^{*}$ up until the last passage time $\lambda_{\alpha^{*}}$. This is due to the higher initial values $Y_{0}$ as compared to $\alpha$, especially for the companies with low CDS spreads. But it is equally true that once the process goes below $\alpha^{*}$, the chance it recovers to $\alpha^{*}$ is small. The last observation can be confirmed by the densities with $c=0$ (in blue) and $c=\frac{2\alpha}{3}$ (in orange): they are almost identical.
Finally, in drawing the graphs in this section, we used the explicit formula in Corollary 4.2 and (4.13) for Figure 1. For Figure 2, we used, in writing the normalized version (with *), $\mathbb{P}_{x^{*}}(\lambda_{\alpha^{*}}\in{\rm d}t)=\frac{p\left(t;x^{*},\alpha^{*}\right)}{s(\alpha^{*})-s(\ell)}{\rm d}t$ where $p(t;x^{*},z^{*})=\frac{1}{2\sqrt{2\pi t}}\exp\left(-M(x^{*}+z^{*})-\frac{M^{2}t}{2}-\frac{(x^{*}-z^{*})^{2}}{2t}\right)$, $s(x^{*})=\frac{1}{2M}(1-e^{-2Mx^{*}})$ and $\ell=$-$\infty$. The initial value $x^{*}$ in the above equation is $\frac{1}{\sigma}\ln(Y_{0})$. In Figure 3, we consider the occupation time of the level above $\alpha^{*}$ before the process $\frac{1}{\sigma}\ln(Y)$ hits some level $\frac{1}{\sigma}\ln(c)=c^{*}<\alpha^{*}$. We denote this time by $\lambda_{\alpha^{*}}^{A,c}$. Its Laplace transform can be easily calculated using (3.1) and is given $\mathbb{E}_{x^{*}}\left[e^{-q\lambda_{\alpha^{*}}^{A,c}}\right]=\frac{\frac{e^{-(\sqrt{M^{2}+2q}+M)x^{*}}}{k}}{\frac{e^{-(\sqrt{M^{2}+2q}+M)\alpha^{*}}}{k}+\frac{1}{2M}(e^{-2Mc^{*}}-e^{-2M\alpha^{*}})}$ with $k=(\sqrt{M^{2}+2q}+M)e^{-(\sqrt{M^{2}+2q}-M)\alpha^{*}}$ as derived in Appendix A.2.
Figure 3. Probability density function (produced by inverting Laplace transform) of the occupancy time above $\alpha^{*}$ before the normalized asset-to-debt ratio hits level $c^{*}$ (starting at $y^{*}=\frac{1}{\sigma}\ln(Y_{0})$) based on the estimated parameters. Three levels of $c$ (stated as legends) are considered for each company: 0, $\frac{2}{3}\alpha$, and $\alpha-0.001$.
@@
A.1. Excursion-theoretic (Point process) view
We can view $\{X_{t},t>\lambda_{\alpha}\}$ as an excursion from $\alpha$ of infinite length. To pursue this point of view, we review some of the excursion theory based on Getoor (1979) and Rogers and Williams (1994b), in particular its Sections VI. 50, 54 and 55.
Let us return to the original process $X$ and consider an excursion from a regular point $\alpha$. Let the local time at $y\in\mathcal{I}$ be $L(t,y),t\geq 0$. Set $L(t)=L(t,\alpha)$ and define the right inverse $\rho(t,\alpha):=\inf\{s:L(s)>t\}$. It is well known that $\rho$ is a right-continuous and nondecreasing subordinator under $\mathbb{P}_{\alpha}$ and hence
$\mathbb{E}_{\alpha}\left[e^{-q\rho(t,\alpha)}\right]=e^{-tg(q)}$ (A.1)
where
$g(q)=m\{\alpha\}q+\int_{(0,\infty]}(1-e^{-qs})\nu({\rm d}s)$ (A.2)
By using the tail distribution
$c(u):=\nu((u,\infty]),$
the Laplace exponent is rewritten as
$g(q)=m\{\alpha\}q+q\int_{0}^{\infty}e^{-qs}c(s){\rm d}s.$ (A.3)
We define
$r_{q}(x,y):=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qt}{\rm d}L(t,y)\right),\quad x,y\in\mathcal{I},q>0.$
Then, by a change of variable and (A.1),
$r_{q}(\alpha,\alpha)=\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-qt}{\rm d}L(t,y)\right)=\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-q\rho(t,\alpha)}{\rm d}t\right)=\frac{1}{g(q)}.$ (A.4)
$U^{q}f(x)=\int_{\mathcal{I}}m({\rm d}y)f(y)\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qs}{\rm d}L(s,y)\right)$
which is also written as
$U^{q}f(x)=\int_{\mathcal{I}}m({\rm d}y)f(y)G_{q}(x,y)$
by using $G_{q}(x,y)=\int_{0}^{\infty}e^{-qt}p(t;x,y){\rm d}t$. See (2.3) for its exact form. By comparison, $G_{q}(x,y)=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qs}{\rm d}L(s,y)\right)$ a.e. $y$. But by the right continuity in $y$, we have the equality for every $y\in\mathcal{I}$. We obtain from (A.4) that $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$ and hence
Remark A.1.
Let $X$ be a linear diffusion with regular point $\alpha\in\mathcal{I}$. We consider the local time at $\alpha$ and excursions from this point. With $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$, (A.3) becomes
$\int_{0}^{\infty}e^{-qs}c(s){\rm d}s=\frac{1}{qG_{q}(\alpha,\alpha)},$
which can be used to obtain $c(u)=\nu((u,\infty])$ through the inverse Laplace transform. This $c(u)$ corresponds to the rate at which the first excursion of length greater than $u$ occurs. To consider an excursion of infinite length, which is closely related to a post-last passage time process discussed in Section 3.2, let us take $q\downarrow 0$ in (A.1) to find
$\mathbb{P}_{\alpha}(\rho(t,\alpha)<\infty)=e^{-t\nu(\{+\infty\})}$
by (A.2). Moreover, from $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$, we have
$\nu(\{+\infty\})=\lim_{q\downarrow 0}\frac{1}{G_{q}(\alpha,\alpha)}:=\frac{1}{G_{0}(\alpha,\alpha)}.$
$\diamondsuit$
A.2. On computations when non-killing (interior) state is made to be killing
Let us comment on the killing boundary involved in the computation of (3.1). Take an example of (I) Brownian motion with negative drift $M$ (as is the normalized $Y$ process in Section 2.3) with two natural boundaries $\ell=$-$\infty$ and $r=+\infty$, of which $\ell$ is a killing boundary. We, of course, use the Green functions $G(\alpha,\alpha)$ and $G_{q}^{A}(x,\alpha)$ for this process to compute the Laplace transform (3.1). Next, let us consider the case (II) where we set a killing boundary, other than natural boundaries, at some $c\in(\ell,r)=\mathcal{I}$. Then, we should use the q-Green function $G_{q}^{A}(x,\alpha)$ for the process “Brownian motion with drift on $(c,+\infty)$ killed at cc”, which is not the same as that for simple Brownian motion with drift.
For case (I), the increasing and decreasing solutions to $\mathcal{G}f=qf$ are given by $\psi_{q}(x)=e^{(\sqrt{M^{2}+2q}-M)x}$ and $\phi_{q}(x)=e^{-(\sqrt{M^{2}+2q}+M)x}$. For case (II), let these solutions be denoted by $\psi_{q}^{c}$ and $\phi_{q}^{c}$ which are linear combinations of $\psi_{q}$ and $\phi_{q}$. For the condition at the killing boundary $c$ we refer the reader to Borodin and Salminen (2002, Chapter II.1. Setion 10). This condition is given by
$\psi_{q}^{c}(c+)=0.$
We have $\psi_{q}^{c}(c)=a_{1}\psi_{q}(c)+a_{2}\phi_{q}(c)$ from which we obtain
$a_{1}=e^{-(\sqrt{M^{2}+2q}+M)c},\quad a_{2}=-e^{(\sqrt{M^{2}+2q}-M)c}.$
As there is no boundary condition for $\phi_{q}^{c}$, we set $\phi_{q}^{c}(x)=\phi_{q}(x)$. Thus, we have
$\psi^{c}_{q}(x)=a_{1}\psi_{q}(x)+a_{2}\phi_{q}(x)=e^{-(\sqrt{M^{2}+2q}+M)c}e^{(\sqrt{M^{2}+2q}-M)x}-e^{(\sqrt{M^{2}+2q}-M)c}e^{-(\sqrt{M^{2}+2q}+M)x}$
$\phi_{q}^{c}(x)=e^{-(\sqrt{M^{2}+2q}+M)x}$
$(\phi_{q}^{c})^{+}(x)=-(\sqrt{M^{2}+2q}+M)e^{-(\sqrt{M^{2}+2q}-M)x}$
together with Wronskian $w_{q}^{c}=2\sqrt{M^{2}+2q}\cdot e^{-(\sqrt{M^{2}+2q}+M)c}$. Then, $G_{q}^{A}(x,\alpha)$ with $x\geq\alpha$ for case (II) is given by
$\frac{-1}{(\phi_{q}^{c})^{+}(\alpha)}\phi_{q}^{c}(x)\left(\frac{-(\phi_{q}^{c})^{+}(\alpha)}{w_{q}^{c}}\psi_{q}^{c}(\alpha)+\frac{(\psi_{q}^{c})^{+}(\alpha)}{w_{q}^{c}}\phi_{q}^{c}(\alpha)\right)=\frac{e^{(\sqrt{M^{2}+2q}-M)\alpha}\cdot e^{-(\sqrt{M^{2}+2q}+M)x}}{(\sqrt{M^{2}+2q}+M)}.$
Finally, as the introduction of the killing boundary $c$ does not change the scale function, case (II) Green function is given by $G^{c}(\alpha,\alpha)=s(\alpha)-s(c)$. This is equal to the Green function of the process in case (I) with $\ell$ just being replaced by $c$. This equality is in fact true for every diffusion, not just Brownian motion with drift. For the details of $G^{A}_{q}(\cdot,\cdot)$, we refer the reader to Egami and Kevkhishvili (2025).
A.3. Proof of Lemma 3.1
@@ A.2. On computations when non-killing (interior) state is made to be killing
A.4. Proof of Proposition 3.3
For the post-last passage time process, we need only to consider a point $y<\alpha$. In view of (3.5), (3.3) becomes
This means that the post-last passage time process is obtained equivalently via the transform $g(\cdot):=s(\alpha)-s(\cdot)$. Call this transform also $X^{h_{\alpha}}$ and its scale function and speed measure $s^{g}$ and $m^{g}$, respectively. By the general result for the transform of this kind (see Appendix A.3), we have
$=\int_{y}^{\alpha}(s(\alpha)-s(\eta))m({\rm d}\eta)<\infty.$
The last finiteness result is due to $\alpha$ being a regular point for $X$. This proves that $\alpha$ is an entrance boundary for the $g$-transform, which is equivalent to the post-last passage time process $X^{h_{\alpha}}$ (see [Karlin and Taylor (1981), Chap.15, Table 6.2]).
A.5. Proof of Proposition 4.1
Using the fact that $J$ is distributed exponentially with rate 1 and independent of $\lambda_{\alpha}$, we have for $T\in\mathbb{R}_{+}$
$\mathbb{P}_{Y_{0}}(\xi\leq T)$ $=\mathbb{P}_{Y_{0}}(\lambda_{\alpha}\leq T-J)=\int_{0}^{T}\mathbb{P}_{Y_{0}}(\lambda_{\alpha}\leq T-x)e^{-x}{\rm d}x$
$=\int_{0}^{T}\mathbb{P}_{Y_{0}}(0<\lambda_{\alpha}\leq T-x)e^{-x}{\rm d}x+\int_{0}^{T}\mathbb{P}_{Y_{0}}(\lambda_{\alpha}=0)e^{-x}{\rm d}x.$
From (2.5) and (4.1), $\lambda_{\alpha}=\sup\{t:\frac{1}{\sigma}\ln(Y_{t})=\alpha^{*}\}$. The last passage time distribution is known: Proposition 4 in Salminen (1984) and Proposition 3.1 in Egami and Kevkhishvili (2020) provide
$\mathbb{P}_{y}(\lambda_{\alpha}\in{\rm d}u)=\frac{p(u;y,\alpha)}{G(\alpha,\alpha)}{\rm d}u\quad\text{and}\quad\mathbb{P}_{y}(\lambda_{\alpha}=0)=1-\frac{G(y,\alpha)}{G(\alpha,\alpha)}$
where $G(\alpha,\alpha)$ is available in (2.1). It can be confirmed that $\mathbb{P}_{Y_{0}}(\lambda_{\alpha}=0)=0$ when $Y_{0}\geq\alpha$ by (2.1) and Assumption 1. Since the transition density $p$ for the process $\frac{1}{\sigma}\ln(Y_{t})$, a Brownian motion with drift $M$, is available, we can evaluate the integral in the above equation explicitly to obtain (4.9).
A.6. Proof of Proposition 4.2
Define $Z^{*}_{t}:=\alpha^{*}-Z_{t}$. Then, by (4.1)
where $(W^{*}_{t})_{t\geq 0}:=(-W_{t})_{t\geq 0}$ is a standard Brownian motion. We have used the fact that $M\coth(-M\cdot x)=M\coth(M\cdot x)$. We have $\tau=\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{u}<\alpha^{*}\}}{\rm d}u=J\right)=\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)$ and
$\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]$ $=\lim_{z^{\prime}\uparrow\alpha^{*}}\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.04861pt\mbox{l}_{\{Z_{u}<\alpha^{*}\}}{\rm d}u=J\right)}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\mid Z_{0}=z^{\prime}\right]$
$=\lim_{z\downarrow 0}\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.04861pt\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z^{*}_{\tau}\geq\alpha^{*}-Q\}}\mid Z_{0}^{*}=z\right].$ (A.6)
By this translation to $Z^{*}$, (4.6) becomes $U(z)=\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.04861pt\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z^{*}_{\tau}\geq\alpha^{*}-Q\}}\mid Z_{0}^{*}=z\right]$ and (4.7) is now
$\frac{1}{2}U^{\prime\prime}(z)+|M|\coth(|M|\cdot z)U^{\prime}(z)-\left(1+\gamma\right)U(z)$ $=-\mbox{1}\hskip-2.5pt\mbox{l}_{[\alpha^{*}-Q,+\infty)}(z),\quad z\in(0,\infty).$ (A.7)
(i) Consider the case when $z\geq\alpha^{*}-Q$. Then (A.7) is reduced to
$\psi_{1}(z)=\frac{\sqrt{2}\sinh\left(|M|b_{1}z\right)}{\sqrt{\pi}b_{1}\sinh(|M|\cdot z)},\quad\phi_{1}(z)=\frac{\sqrt{\pi}e^{-|M|b_{1}z}}{\sqrt{2}\sinh(|M|\cdot z)},$ (A.8)
(ii) Let $0<z<\alpha^{*}-Q$. In this case, (A.7) is reduced to
As in the case (i) above, its fundamental solutions are given by $\psi_{1}(z)$ and $\phi_{1}(z)$ in (A.8). Therefore, in this case we have $U(z)=d\psi_{1}(z)$ with some constant $d$. We have eliminated the second fundamental solution because $U(z)$ is bounded when $z\to 0$.
Finally, $\lim_{z\downarrow 0}U(z)=\lim_{z\downarrow 0}d\cdot\psi_{1}(z)=d\cdot\sqrt{\frac{2}{\pi}}$, which proves (4.10) in view of (A.6).
A.6.1. Proof of Corollary 4.1
$\mathbb{P}_{\alpha^{*}}\left(Z_{\tau}\leq Q\right)=\lim_{\gamma\downarrow 0}\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right].$
Thus, we obtain the desired result by taking the limit $\gamma\downarrow 0$ in (4.10).
A.6.2. Proof of Corollary 4.2
Due to $Y_{\xi}\overset{d}{\sim}e^{\sigma Z_{\tau}}$, we have $K^{B}(\xi)\overset{d}{\sim}1-e^{\sigma Z_{\tau}}$ by (4.5). The density function of $Z_{\tau}$ is obtained by differentiating (4.11) with respect to $Q$:
Then, the density of the transform $1-e^{\sigma Z_{\tau}}$ is given by
Finally, $\int_{1-e^{\sigma\alpha^{*}}}^{1}x\mathbb{P}_{\alpha^{*}}(K^{B}(\xi)\in{\rm d}x)$ provides (4.12).
A.7. Detailed comments on Step 1 in Section 4.2
@@
We estimate the parameters of the asset process in (2.5) using the data of the firm’s equity $E$ and debt $B$ available up to the current time. The debt process $B$ represents a certain amount of debt to be repaid and we let it be the sum of short-term debt and a half of long-term debt (as in Moody’s KMV approach) since empirical evidence has demonstrated that such level is an appropriate default threshold for the firm’s assets.
We use the option-theoretic approach of Merton (1974) where the value of equity represents the value of the European call option written on the firm’s assets with the strike price equal to the future value of debt. We estimate the asset parameters using the maximum likelihood method of Duan (1994), Duan (2000), and Lehar (2005). We assume that debt $B$ grows at the risk-free rate $r$. Assuming that debt $B$ is exogenous, the log-likelihood function to be maximized is given by
@@
Acknowledgment
We thank Tomohiro Koike (Graduate School of Economics, Kyoto University) for his valuable comments regarding Proposition 3.2.